OpenStudy (anonymous):

The average velocity of a car over a certain time interval is 37 mi/h. If the velocity of the car was 65 mi/h at the end of this interval, what was its initial velocity. Assume that acceleration was constant.

OpenStudy (anonymous):

you are not given the time interval?

OpenStudy (anonymous):

i guess we could call it \(t\) and the average velocity is the slope (since acceleration is constant) then solve \[\frac{65-x}{t}=37\] for \(x\)

OpenStudy (loser66):

is it not that average velocity \(v=\dfrac{v_{final}+v_0}{2}\) we have \(v_{final}\)= 65, and \(v= 37\) we can count the \(v_0\) , right?

OpenStudy (anonymous):

I'm working on it and found 8 mi/h ?? am i correct:d

OpenStudy (loser66):

I don't know why, I got 9mi/h. My work is 37 = \(\dfrac{65 +v_0}{2}\)\(\rightarrow\) 74 = 65 +\(v_0\)\(\rightarrow\) \(v_0\) = 74-65 =9

OpenStudy (anonymous):

yep you're right.d

OpenStudy (anonymous):

thanks....

OpenStudy (loser66):

ok